# Teaching Spell Checkers to Teach: Pedagogical Program Synthesis for Interactive Learning

### v1

@@ 2.3. AI Pedagogical Techniques
AI-assisted literacy support has evolved from error correction systems (contextual_gamon_2008) to more sophisticated instructional approaches that adapt to individual learners. The emergence of large language models has particularly enabled generative methods that create personalized learning experiences. Goodman et al. (lampost_goodman_2022) designed LaMPost, an AI-assisted writing tool for adults with dyslexia that helps compose emails while preserving user agency. Wang et al. (10.1145/3038535.3038542) developed conceptual motivation models specifically for students with dyslexia, recognizing that practical assistive learning tools must take into account motivational factors alongside functional capabilities. Leong et al. (putting_leong_2024) demonstrated that AI-enabled context personalization for vocabulary learning improves motivation by situating words within personally relevant scenarios. Attygalle et al. (10.1145/3708359.3712073) explored text-to-image generation for vocabulary learning using the keyword method, finding that externalizing memorable visual links significantly increases memory retention. Peng et al. (storyfier_peng_2023) developed Storyfier, exploring vocabulary support through narrative contexts generated by text models. Wang et al. (learnmate_wang_2025) extended this with LearnMate, providing personalized learning plans and adaptive support. Investigation of AI-based support in speech-language pathology has documented both opportunities and concerns about bias and clinical validity (exploring_lewis_2025), with particular attention to providing equitable care for culturally and linguistically diverse children. Design explorations of AI-supported home practice in speech therapy reveal how varying levels of AI oversight can provide informational, emotional, and practical support to parents implementing therapy techniques (10.1145/3706598.3713986), although questions remain about alignment with professional therapeutic goals.
@@
Conversational and Socratic approaches have emerged for fostering critical thinking through interactive dialogue. Ruan et al. (10.1145/3397481.3450648) developed EnglishBot, an AI-powered conversational system for second language learning that adapts feedback based on student interaction, demonstrating greater engagement compared to traditional listen-and-repeat interfaces. Systems using AI to probe student understanding through targeted questioning enable scalable oral assessment (socratic_hung_2024), while question-driven exploration for personalized education (socratiq_jabbour_2025) and investigations of whether AI partners can empower learners to ask critical questions (ai_maiti_2025) suggest that appropriately designed systems scaffold metacognitive skills. Metacognitive tutoring agents have proven to be effective in teaching inquiry-driven modeling in science education, demonstrating that intelligent tutoring systems can scaffold authentic scientific inquiry processes (10.1145/2678025.2701398). However, these conversational approaches, while effective for verbal reasoning, may not optimally support the hands-on, constructive activities central to structured word inquiry.
@@
Program synthesis offers an approach particularly well suited to inquiry-based learning contexts requiring structured artifacts. Unlike conversational AI that responds to queries, program synthesis can generate interactive exercises and examples tailored to specific learning objectives. Theoretical foundations for program synthesis (program_gulwani_2017) have been extended by demonstrations that large language models effectively perform synthesis tasks (program_austin_2021), including interactive synthesis by augmented examples (interactive_zhang_2020). Recent work has explored how hypothesis search enables inductive reasoning with language models (wang2024hypothesis), providing mechanisms for iterative refinement of generated artifacts. Synthesis and verification pipelines demonstrate how LLM-generated outputs can be iteratively corrected through automated verification and feedback loops (10.1145/3731209), achieving high success rates through repeated refinement cycles. Kazemitabaar et al. (codeaid_kazemitabaar_2024) evaluated CodeAid in classroom deployment, documenting how LLM-based synthesis balances exploration with scaffolding. For inquiry-based spelling instruction, program synthesis enables dynamic generation of morphological investigations adapted to learners’ current hypotheses, providing the structured yet flexible framework necessary for effective SWI implementation at scale.
@@ 3. User Experience
At a high level, SPIRE transforms traditional spell checking into an interactive process of linguistic inquiry. When a learner misspells a word, SPIRE invites them to explore its structure, meaning, and sound rather than simply replacing it with a correction. Each spelling mistake becomes an opportunity to reason about English orthography through short, guided activities embedded directly in the writing interface.
@@ 3.1. Usage Scenario
Matilda opens her word processor, which has SPIRE integrated into it. As she types, she writes: “I like how the art of constractd.” The system flags constractd as a possible error by enclosing it in a red rectangle. When Matilda clicks on the word, the red highlight transitions into a red text indicator, and an inquiry canvas opens below her writing (Fig. 2a1).
The system begins by asking Matilda to use the word in a sentence—a diagnostic step to determine whether she understands its meaning. She clicks on a microphone icon (Fig. 2a1) and says aloud: “The builder constructed the building.” The system transcribes her response and verifies that her usage reflects an appropriate understanding of construct. Since her sense of meaning is correct, it affirms her answer (Fig. 2a3) and proceeds to the next inquiry.
Next, SPIRE invites Matilda to “box the base“ of the word (Fig. 2b1), a structured word inquiry practice where learners identify the morphemic base. Using a paint roller tool, Matilda drags across the portion she believes to be the base and highlights stract (Fig. 2b2). Although not fully correct, this action shows that she recognizes the structural pattern embedded within construct. The system records this as partial evidence of her morphological awareness and proceeds to test a new hypothesis related to graphemes.
@@
Finally, SPIRE prompts Matilda to suggest words that share the same base <struct> (Fig. 2c1). She types structure and insstruct (Fig. 2c2). The system affirms structure but flags insstruct for having an extra <s>, explaining the morphological constraint (Fig. 2c3). This final step reinforces her understanding of morphological families and related word forms before she returns to her essay.
Through these brief, multimodal interactions–speaking, dragging, typing—Matilda not only corrects her mistake but also deepens her understanding of word structure and related forms. Each step is conversational, adaptive, and pedagogically grounded, embodying how SPIRE turns everyday spelling errors into opportunities for structured linguistic exploration.
@@ 4. Pedagogical Program Language
To develop SPIRE, we need to encode the dynamic conversational practice of SLPs into something a computer can assemble in real time from a single misspelling and its context. Our pedagogical programming language (i.e., DSL) formalizes how SLPs reason, question, and guide discovery during therapy as composable program primitives that can be synthesized into concise, adaptive instructional sequences. Crucially, our DSL is grounded in practitioner concepts and built with practitioners. We first inventoried canonical SWI moves and terminology from established texts (books written by expert practitioners, instructional material used in graduate programs, and skill development workshops (hegland2021beneath; barnett2024vocabulary; bowers2006teaching; constantine2024structured; loveless2024high)) into a spreadsheet. Next, working with a practicing SLP, who is also a co-author, we organized the moves as program primitives, including language primitives, hypothesis templates, action primitives, and control and compositional primitives. For this collaborative work, we utilized digital whiteboards to organize the concepts, created sketches of program logic and instructional moves, and also produced storyboards of envisioned learning episodes. This was an iterative process that involved weekly virtual meetings with the research team, spanning over six months. During this process, we also consulted with two other practicing SLPs who have over 10 years of experience to receive feedback on the evolving DSL.
@@ 4.1. Language Knowledge Primitives
morphemes Meaning-bearing parts composing the word; focus of “box the base” and word-sum construction. ["reach", "-ing"]
bases Core lexical root carrying meaning across a family; anchors family-based pattern recognition. ["reach"]
prefixes Bound morphemes before the base; invite prefix-meaning inquiries and family extension. none
suffixes Bound morphemes following the base; used in suffixing-convention and inflectional change prompts. ["-ing"]
graphemes Letter or letter-team units representing phonemes; supports grapheme identification and comparison. ["r", "ea", "ch", "i", "ng"]
[…]
related_words Morphological or orthographic relatives for matrices and “in/out” sorting; surface consistent base spelling. “reach”, “reaching”, “overreach”; contrasts: “teach”, “preach”
etymology Historical origin explaining modern spelling stability; supports etymological reasoning. Old English reccan (“to stretch out”)
homophones Same sound, different spelling or meaning; prompts meaning-based distinction. none
semantic_appropriateness Ensures the written word matches intended meaning; drives meaning-first inquiry when mismatched. Intended “extending or stretching out” (“She is reaching for the shelf.”)
syntactic_correctness Checks grammatical fit; distinguishes morphological from typographic errors. Progressive verb fits: “She is reaching…”
meaning_understood If unclear, prioritizes definition/denotation prompts before structure/GPC work. Learner typed “reeching” intending “reaching”.
context_sentence Provides authentic sentence evidence; used in meaning or boundary reasoning. “She was reeching for the book.”
@@ 4.2. Hypotheses Template Primitives
Our DSL formalizes the inquiry logic of structured word inquiry as a set of reusable hypothesis templates. Each template represents a recurring pattern of reasoning that enables the learner to test and refine a hypothesis about how a word works. It expresses a single pedagogical move typical in SWI as a structured quintuple consisting of preconditions, evidential features, action base, warrant, and learning effect. Preconditions state when a line of inquiry is warranted. They are conditional predicates over an observable situation, including the type of deviation or error in a learner’s spelling production, the linguistic characteristics of the intended word, and basic usage context, such as semantic appropriateness. Concretely, they are represented as a combination of interpretable Boolean predicates computed over the linguistic primitives as well as diagnostic descriptors in natural language. The predicate provides a formal check for whether the template applies, while the descriptor captures practitioner judgment that is hard to encode as a formal rule.
For instance, in the grapheme identification hypothesis template (H8)(Table. 2), the symbolic guard checks that the that the deviation is grapheme–local rather than structural: morphology and boundary errors are not implicated ($\neg$(prefix_error $\vee$ suffix_error $\vee$ segmentation_error)), no suffixing convention applies ($\neg$ suffixing_change_applies), the phoneme sequence of the attempt essentially matches the target (phoneme_match $\leq\varepsilon$), the mismatch is small and local (grapheme_mismatch_count $\in\{1,2\}$), and morpheme boundaries are preserved (morpheme_boundaries_preserved). The paired diagnostic descriptor states this in practitioner terms e.g., “sounds correct but uses a nonstandard letter team.” Thus reechreach satisfies the guard (phonology intact, one letter–team substitution), whereas runingrunning is excluded by the suffixing-change check and alota lot by the segmentation predicate. At runtime, an LLM can assign soft confidence to descriptors, allowing preconditions to serve as both interpretable constraints and probabilistic cues for selecting the most pedagogically relevant hypothesis.
@@ 4.3. Compositional Primitives
[…]
gray!50
[…]
H1     Meaning Clarification     Meaning unclear or semantically inappropriate.     Define intended meaning; contextual prompt.     Meaning aligned; enables next inquiry.
H2     Base Denotation     Base + affix present; meaning known.     Identify main meaning part (base).     Semantic core anchored in base.
H3     Morphological Structure     Affix $\geq$1; single base.     “Box the base” or decompose parts.     Base verified; structure understood.
H4     Word-Sum Construction     Decomposable into morphemes.     Build word step-by-step (synthetic sum).     Morphemes combined → correct form.
H5     Spelling Convention     Suffixing rule applies (e.g., doubling, e-drop).     Inspect suffix rule via word-sum.     Orthographic rule induced.
H6     Word Matrix     Base stable across $\geq$2 relatives.     Assemble word matrix grid.     Family pattern generalized.
H7     Word Boundaries     Segmentation or spacing error.     Segment aloud, retype spaced form.     Correct word boundaries restored.
H8     Grapheme Identification     Small grapheme mismatch ($\leq$2).     Identify graphemes, compare to target.     Grapheme–phoneme alignment achieved.
H9     Etymology Inquiry     Historical root available.     Trace origin; discuss spelling stability.     Origin-based spelling rationale.
H10     Morphological Relatives     Base shared by other words.     Sort IN/OUT relatives; construct family.     Base spelling reinforced by family.
H11     Morpheme Verification     Base/affix uncertainty.     Verify morphemes across known words.     Morpheme consistency confirmed.
H12     Etymological Relatives     Common ancestor offers insight.     Compare historical cousins.     Etymological family explained.
H13     False Relatives     Optional; visual similarity only.     Contrast look-alike vs. true relatives.     False relatives excluded.
H14     Phoneme–Grapheme Mapping     Phonology $\approx$ target.     Map sounds to graphemes (phoneme table).     Sound–spelling map aligned.
H15     Pronunciation Variation     Family shows sound shifts.     Compare relatives with shared grapheme.     Stable spelling despite sound change.
H16     GPC Pattern Consistency     Consistent spelling, varied pronunciation.     Compare family spellings.     Pattern consistency generalized.
H17     Homophone Distinction     $\geq$1 homophone exists.     Sort by meaning; contrast forms.     Correct form tied to meaning.
H18     Similar Word Comparison     Similar orthography detected.     Visual contrast task.     Subtle form difference noticed.
Table 2. This table consists of the abbreviated summary of SPIRE Hypothesis Templates.
@@ 5. System Architecture
This architecture addresses a fundamental limitation in existing spelling tools. Conventional spell checkers detect errors and suggest corrections based on string similarity, treating spelling as a mechanical transcription problem. SPIRE must instead infer why a misspelling occurred, what it reveals about the learner’s linguistic reasoning, and which line of inquiry will advance their orthographic understanding. The spelling inquiry engine, implemented in Python, orchestrates this process through five modules: a spell checker identifies target words from context, a word property synthesizer extracts morphological and phonological features, an error analyzer diagnoses likely causes, a spelling inquiry engine composes hypothesis templates into executable programs following the DSL notation, and an interface generator renders these programs as interactive components. Together, these modules transform the DSL’s declarative specifications into adaptive, just-in-time instruction tailored to individual learner errors.

### v2

@@ 2.3. AI Pedagogical Techniques
AI-assisted literacy support has evolved from error correction systems (Gamon et al., 2008) to more sophisticated instructional approaches that adapt to individual learners. The emergence of large language models has particularly enabled generative methods that create personalized learning experiences. Goodman et al. (Goodman et al., 2022) designed LaMPost, an AI-assisted writing tool for adults with dyslexia that helps compose emails while preserving user agency. Wang et al. (Wang et al., 2017) developed conceptual motivation models specifically for students with dyslexia, recognizing that practical assistive learning tools must take into account motivational factors alongside functional capabilities. Leong et al. (Leong et al., 2024) demonstrated that AI-enabled context personalization for vocabulary learning improves motivation by situating words within personally relevant scenarios. Attygalle et al. (Attygalle et al., 2025) explored text-to-image generation for vocabulary learning using the keyword method, finding that externalizing memorable visual links significantly increases memory retention. Peng et al. (Peng et al., 2023) developed Storyfier, exploring vocabulary support through narrative contexts generated by text models. Wang et al. (Wang et al., 2025) extended this with LearnMate, providing personalized learning plans and adaptive support. Investigation of AI-based support in speech-language pathology has documented both opportunities and concerns about bias and clinical validity (Lewis et al., 2025), with particular attention to providing equitable care for culturally and linguistically diverse children. Design explorations of AI-supported home practice in speech therapy reveal how varying levels of AI oversight can provide informational, emotional, and practical support to parents implementing therapy techniques (Dangol et al., 2025), although questions remain about alignment with professional therapeutic goals.
@@
Conversational and Socratic approaches have emerged for fostering critical thinking through interactive dialogue. Ruan et al. (Ruan et al., 2021) developed EnglishBot, an AI-powered conversational system for second language learning that adapts feedback based on student interaction, demonstrating greater engagement compared to traditional listen-and-repeat interfaces. Systems using AI to probe student understanding through targeted questioning enable scalable oral assessment (Hung et al., 2024), while question-driven exploration for personalized education (Jabbour et al., 2025) and investigations of whether AI partners can empower learners to ask critical questions (Maiti and Goel, 2025) suggest that appropriately designed systems scaffold metacognitive skills. Metacognitive tutoring agents have proven to be effective in teaching inquiry-driven modeling in science education, demonstrating that intelligent tutoring systems can scaffold authentic scientific inquiry processes (Joyner and Goel, 2015). However, these conversational approaches, while effective for verbal reasoning, may not optimally support the hands-on, constructive activities central to structured word inquiry.
@@
Program synthesis offers an approach particularly well suited to inquiry-based learning contexts requiring structured artifacts. Unlike conversational AI that responds to queries, program synthesis can generate interactive exercises and examples tailored to specific learning objectives. Theoretical foundations for program synthesis (Gulwani et al., 2017) have been extended by demonstrations that large language models effectively perform synthesis tasks (Austin et al., 2021), including interactive synthesis by augmented examples (Zhang et al., 2020). Recent work has explored how hypothesis search enables inductive reasoning with language models (Wang et al., 2024), providing mechanisms for iterative refinement of generated artifacts. Synthesis and verification pipelines demonstrate how LLM-generated outputs can be iteratively corrected through automated verification and feedback loops (Ma and Agrawala, 2025), achieving high success rates through repeated refinement cycles. Kazemitabaar et al. (Kazemitabaar et al., 2024) evaluated CodeAid in classroom deployment, documenting how LLM-based synthesis balances exploration with scaffolding. For inquiry-based spelling instruction, program synthesis enables dynamic generation of morphological investigations adapted to learners’ current hypotheses, providing the structured yet flexible framework necessary for effective SWI implementation at scale.
@@ 3. User Experience
At a high level, SPIRE transforms traditional spell checking into an interactive process of linguistic inquiry. When a learner misspells a word, SPIRE invites them to explore its structure, meaning, and sound rather than simply replacing it with a correction. Each spelling mistake becomes an opportunity to reason about English orthography through short, guided activities embedded into the writing interface.
@@ 3.1. Usage Scenario
Matilda opens her word processor, which has SPIRE integrated into it. As she types, she writes: “I like how the art of constractd.” The system flags constractd as a possible error by enclosing it in a red rectangle. When Matilda clicks on the word, the red highlight transitions into a red text indicator, and an inquiry canvas opens below her writing (Fig. 2A1).
The system begins by asking Matilda to use the word in a sentence—a diagnostic step to determine whether she understands its meaning. She clicks on a microphone icon (Fig. 2A1) and says aloud: “The builder constructed the building.” The system transcribes her response and verifies that her usage reflects an appropriate understanding of construct. Since her sense of meaning is correct, it affirms her answer (Fig. 2A3) and proceeds to the next inquiry.
Next, SPIRE invites Matilda to “box the base“ of the word (Fig. 2B1), a structured word inquiry practice where learners identify the morphemic base. Using a paint roller tool, Matilda drags across the portion she believes to be the base and highlights stract (Fig. 2B2). Although not fully correct, this action shows that she recognizes the structural pattern embedded within construct. The system records this as partial evidence of her morphological awareness and proceeds to test a new hypothesis related to graphemes.
@@
Finally, SPIRE prompts Matilda to suggest words that share the same base <struct> (Fig. 2C1). She types structure and insstruct (Fig. 2C2). The system affirms structure but flags insstruct for having an extra <s>, explaining the morphological constraint (Fig. 2C3). This final step reinforces her understanding of morphological families and related word forms before she returns to her essay.
Through these brief, multimodal interactions–speaking, dragging, typing—Matilda not only corrects her mistake but also deepens her understanding of word structure and related forms. By condensing inquiry into a short, bounded sequence, the system allows her to engage in a "micro-lesson" without significantly derailing her drafting process. This interaction model illustrates how SPIRE seeks to balance writing fluency with learning, turning selected spelling errors into manageable opportunities for structured linguistic exploration.
@@ 4. Pedagogical Program Language
To develop SPIRE, we need to encode the dynamic conversational practice of SLPs into something a computer can assemble in real time from a single misspelling and its context. Our pedagogical programming language (i.e., DSL) formalizes how SLPs reason, question, and guide discovery during therapy as composable program primitives that can be synthesized into concise, adaptive instructional sequences. Crucially, our DSL is grounded in practitioner concepts and built with practitioners. We first inventoried canonical SWI moves and terminology from established texts (books written by expert practitioners, instructional material used in graduate programs, and skill development workshops (Hegland, 2021; Barnett and Squires, 2024; Bowers, 2006; Constantine and Lee, 2024; Loveless and Hamilton, 2024)) into a spreadsheet. Next, working with a practicing SLP, who is also a co-author, we organized the moves as program primitives, including language primitives, hypothesis templates, action primitives, and control and compositional primitives. For this collaborative work, we utilized digital whiteboards to organize the concepts, created sketches of program logic and instructional moves, and also produced storyboards of envisioned learning episodes. This was an iterative process that involved weekly virtual meetings with the research team, spanning over six months. During this process, we also consulted with two other practicing SLPs who have over 10 years of experience to receive feedback on the evolving DSL.
@@ 4.2. Hypotheses Template Primitives
Our DSL formalizes the inquiry logic of structured word inquiry as a set of reusable hypothesis templates. Each template represents a recurring pattern of reasoning that enables the learner to test and refine a hypothesis about how a word works. It expresses a single pedagogical move typical in SWI as a structured quintuple consisting of preconditions, evidential features, action base, warrant, and learning effect. Preconditions state when a line of inquiry is warranted. They are conditional predicates over an observable situation, including the type of deviation or error in a learner’s spelling production, the linguistic characteristics of the intended word, and basic usage context, such as semantic appropriateness. Concretely, they are represented as a combination of Boolean predicates computed over the linguistic primitives as well as diagnostic descriptors in natural language. The predicate provides a formal check for whether the template applies, while the descriptor captures practitioner judgment that is hard to encode as a formal rule.
For instance, in the grapheme identification template (H8 in Table 2), the symbolic guard checks that the that the deviation is grapheme–local rather than structural: morphology and boundary errors are not implicated ($\neg$(prefix_error $\vee$ suffix_error $\vee$ segmentation_error)), no suffixing convention applies ($\neg$ suffixing_change_applies), the phoneme sequence of the attempt essentially matches the target (phoneme_match $\leq\varepsilon$), the mismatch is small and local (grapheme_mismatch_count $\in\{1,2\}$), and morpheme boundaries are preserved (morpheme_boundaries_preserved). The paired diagnostic descriptor states this in practitioner terms e.g., “sounds correct but uses a nonstandard letter team.” Thus reechreach satisfies the guard (phonology intact, one letter–team substitution), whereas runingrunning is excluded by the suffixing-change check and alota lot by the segmentation predicate. At runtime, an LLM assigns soft confidence to descriptors, allowing preconditions to serve as interpretable constraints and probabilistic cues for selecting the most pedagogically relevant hypothesis.
@@ 4.3. Compositional Primitives
ID Template Key Preconditions Action Base Learning Effect H1 Meaning Clarification Meaning unclear or semantically inappropriate. Define intended meaning; contextual prompt. Meaning aligned; enables next inquiry. H2 Base Denotation Base and affix present; meaning known. Identify main meaning part. Semantic core anchored in base. H3 Morphological Structure 1 or more affixes; single base. “Box the base” or decompose parts. Base verified; structure understood. H4 Word-Sum Construction Decomposable into morphemes. Build word step-by-step. Morphemes combined $\rightarrow$ correct form. H5 Spelling Convention Suffixing rule applies (e.g., doubling, e-drop). Inspect suffix rule via word-sum. Orthographic rule induced. H6 Word Matrix Base stable across 2 or more relatives. Assemble word matrix grid. Family pattern generalized. H7 Word Boundaries Segmentation or spacing error. Segment aloud, retype spaced form. Correct word boundaries restored. H8 Grapheme Identification 2 or less grapheme mismatch. Identify graphemes, compare to target. Grapheme–phoneme alignment achieved. H9 Etymology Inquiry Historical root available. Trace origin; discuss spelling stability. Origin-based spelling rationale. H10 Morphological Relatives Base shared by other words. Sort IN/OUT relatives; construct family. Base spelling reinforced by family. H11 Morpheme Verification Base/affix uncertainty. Verify morphemes across known words. Morpheme consistency confirmed. H12 Etymological Relatives Common ancestor offers insight. Compare historical cousins. Etymological family explained. H13 False Relatives Only visual similarity. Contrast look-alike vs. true relatives. False relatives excluded. H14 Phoneme–Grapheme Map Phonology is similar to target. Map sounds to graphemes. Sound–spelling map aligned. H15 Pronunciation Variation Family shows sound shifts. Compare relatives with shared grapheme. Stable spelling despite sound change. H16 GPC Pattern Consistency Consistent spelling, varied pronunciation. Compare family spellings. Pattern consistency generalized. H17 Homophone Distinction 1 or more homophones exist. Sort by meaning; contrast forms. Correct form tied to meaning. H18 Similar Word Comparison Similar orthography detected. Visual contrast task. Subtle form difference noticed.
Table 2. Summarized Hypothesis Templates of SPIRE.
@@ 5. System Architecture
This architecture addresses a fundamental limitation in existing spelling tools. Conventional spell checkers detect errors and suggest corrections based on string similarity (Kukich, 1992), treating spelling as a mechanical transcription problem. SPIRE must instead infer why a misspelling occurred, what it reveals about the learner’s linguistic reasoning, and which line of inquiry will advance their orthographic understanding. The spelling inquiry engine, implemented in Python, orchestrates this process through five modules: a spell checker identifies target words from context (Fig. 3, step 1), a word property synthesizer extracts morphological and phonological features (Fig. 3, step 2), an error analyzer diagnoses likely causes (Fig. 3, step 3), a spelling inquiry engine composes hypothesis templates into executable programs following the DSL notation (Fig. 3, step 4), and an interface generator renders these programs as interactive components. Together, these modules transform the DSL’s declarative specifications into adaptive, just-in-time instruction tailored to individual learner errors.